OpenStudy (gabbyalicorn):

A circular plate has a diameter of 6 inches. What is the exact area of the plate? A. 36π square inches B. 6π square inches C. 9π square inches D. 3π square inches

OpenStudy (gabbyalicorn):

I think A

OpenStudy (owlcoffee):

Well, you can calculate the area of a circle with the following formula: \[A _{\circ}= \pi r^2\] So, therefore: \[A_ \circ = \pi (6)^2\]

OpenStudy (gabbyalicorn):

I got 113.04

OpenStudy (michele_laino):

the requested area is given by this formula: \[A = \pi \times \frac{{{D^2}}}{4} = \pi \times \frac{{36}}{4} = ...inche{s^2}\]

OpenStudy (owlcoffee):

oooh, it's the diameter, read wrong :(

OpenStudy (gabbyalicorn):

wait so 3.14 x 36 div 4 ?

OpenStudy (michele_laino):

you don't have to multiply by 3.14, you have to write pi only

OpenStudy (gabbyalicorn):

but pi = 3.14

OpenStudy (michele_laino):

yes! nevertheless you have to write pi instead of 3.14

OpenStudy (gabbyalicorn):

okay.

OpenStudy (michele_laino):

so, what dou you get?

OpenStudy (gabbyalicorn):

28.26

OpenStudy (michele_laino):

I got: \[A = 9 \times \pi {\kern 1pt} inche{s^2}\]

OpenStudy (gabbyalicorn):

oh. well it still equals the same thing.

OpenStudy (gabbyalicorn):

brb

OpenStudy (michele_laino):

you have to divide 36 by 4 only

OpenStudy (gabbyalicorn):

which is 9

OpenStudy (gabbyalicorn):

Oh! C :D lol Yay! Thank you! ^u^

OpenStudy (michele_laino):

:)